### v1

# Heterogeneous Mixture of Experts for Remote Sensing Image Super-Resolution

Abstract
Remote sensing image super-resolution (SR) aims to reconstruct high-resolution remote sensing images from low-resolution inputs, thereby addressing limitations imposed by sensors and imaging conditions. However, the inherent characteristics of remote sensing images, including diverse ground object types and complex details, pose significant challenges to achieving high-quality reconstruction. Existing methods typically employ a uniform structure to process various types of ground objects without distinction, making it difficult to adapt to the complex characteristics of remote sensing images. To address this issue, we introduce a Mixture of Experts (MoE) model and design a set of heterogeneous experts. These experts are organized into multiple expert groups, where experts within each group are homogeneous while being heterogeneous across groups. This design ensures that specialized activation parameters can be employed to handle the diverse and intricate details of ground objects effectively. To better accommodate the heterogeneous experts, we propose a multi-level feature aggregation strategy to guide the routing process. Additionally, we develop a dual-routing mechanism to adaptively select the optimal expert for each pixel. Experiments conducted on the UCMerced and AID datasets demonstrate that our proposed method achieves superior SR reconstruction accuracy compared to state-of-the-art methods. The code will be available at https://github.com/Mr-Bamboo/MFG-HMoE.
Index Terms:
Remote sensing images, super-resolution, mixture of experts, upsample, multi-level feature
I Introduction
High-resolution remote sensing imagery provides detailed representations of ground objects, serving critical functions in applications ranging from land-use monitoring [1] to post-disaster assessment [2] and ecological conservation. However, various constraints in imaging conditions and sensor hardware often result in the acquisition of low-resolution images [3], which significantly degrades their effectiveness in downstream applications. Super-resolution (SR) provides a non-physical solution to these resolution limitations by enabling the reconstruction of high-resolution (HR) images from low-resolution (LR) inputs [4].
Early super-resolution (SR) methods were primarily based on interpolation or optimization techniques[5, 6]. However, these approaches often struggle to model the complex nonlinear features present in remote sensing images. In recent years, deep learning-based SR methods have emerged as the dominant approach [7, 8, 9, 10], leveraging carefully designed deep neural networks to reconstruct images with fine structural details. Initially, convolutional neural networks (CNNs) were employed for SR reconstruction. Lideg et al. [11] introduced a CNN with residual connections, which was later modified by Lim et al [12]. Subsequently, attention-based methods gained widespread attention due to their superior long-range modeling capabilities. Zhang et al. proposed RCAN [13], which integrates convolutional operations with channel attention. Lei et al. developed a specialized self-attention mechanism to exploit self-similarity in remote sensing images. Chen et al. introduced HAT[9], which combines multiple attention mechanisms to extract more comprehensive features. Building on HAT, Xiao et al. [10] further incorporated token selection into the feature extraction process to reduce redundant information.
However, current methods primarily rely on a single and fixed structure to process the intricate details of ground objects. This approach struggles to distinguish and reconstruct different types of ground objects effectively, often leading to suboptimal results. To address this limitation, we introduce a Mixture of Experts (MoE) model, where different experts are employed for upsampling.
To better accommodate the complexity of ground objects in remote sensing images, we construct a set of heterogeneous experts. Specifically, these experts are grouped into multiple expert sets, where experts within the same group share the same structure, while those across groups have different architectures. This design allows for specialized upsampling tailored to different types of ground objects. Furthermore, to adapt to this heterogeneous expert design, we propose two key components. First, we introduce a Multi-level Feature Aggregation (MFA) strategy, which aggregates multi-level features from the backbone network to estimate expert activation probabilities, thereby guiding MoE routing. Second, we design a dual-routing mechanism, where an initial routing step selects the most suitable expert group, followed by a second routing step that determines the optimal expert within the selected group. Together, these components form the proposed Multi-level Feature Guided Heterogeneous Mixture of Experts (MFG-HMoE).
The contributions of our work are summarized as follows:
1) We introduce MoEs into the remote sensing image SR task and propose MFG-HMoE, which enables adaptive high-quality reconstruction for each pixel.
2) To handle the complexity of remote sensing scenes, we design a set of heterogeneous experts and propose the MFA strategy along with the dual-routing mechanism to route ground object features to the most suitable expert for processing.
3) Our method outperforms contemporary approaches on the UCMerced [14] and AID [15] datasets, demonstrating its advanced capabilities.
The remainder of this paper is structured as follows: Section II presents a detailed description of our methodology; Section III discusses the experimental evaluation of the super-resolution results; and Section IV concludes the paper.
II Methodology
Figure 1: The flowchart of the proposed MFG-HMoE. The feature extraction network is constructed by stacking RHAGs [9]. The MFA module aggregates the multi-level features from the feature extraction network and feeds them into the two routers. DR-HMoE routes each feature pixel output from the feature extraction network to the optimal expert for processing.
II-A Overview
We define the following components of the super-resolution process: $\mathbf{I}_{LR}$ represents the input low-resolution image, $\mathbf{I}_{SR}$ denotes the output high-resolution image, $h_{\theta}(\cdot)$ is the feature extraction backbone, and $h_{up}(\cdot)$ represents the upsampling head. The super-resolution process can then be formulated as:
$\mathbf{I}_{SR}=h_{up}(h_{\theta}(\mathbf{I}_{LR})).$ (1)
We employ stacked Residual Hybrid Attention Groups (RHAG) [9] as the feature extraction backbone. To enhance the upsampling process, we introduce a Multi-Level Feature Aggregation (MFA) module that consolidates features from multiple levels to activate different experts. Within the upsampling process, we propose a Dual-Routing Heterogeneous Mixture of Experts (DR-HMoE) to enable adaptive pixel-level semantic expression, addressing the complex surface composition characteristics and diverse frequency patterns inherent in remote sensing imagery. Specifically, we design $N$ groups of upsampling experts, where experts within each group share identical structures while maintaining heterogeneity across groups. The expert activation follows a cascading process: first, the optimal expert group is selected based on MFA information, followed by the selection of the most suitable expert within that group using both MFA information and expert-specific cues. The detailed architecture is illustrated in Fig. 1.
II-B Multi-level Feature Aggregation
The Multi-level Feature Aggregation (MFA) module is designed to encapsulate backbone features and provide cues for expert activation. Given $\mathbf{X}^{i}\in\mathbb{R}^{C\times H\times W}$ as the output features from the $i$-th RHAG block, we obtain intermediate hidden features through:
${\mathbf{X}^{i}}^{\prime}=\text{LeakyReLU}(\text{Conv2D}(\mathbf{X}^{i}))$ (2)
where Conv2D and LeakyReLU denote the 2D convolution operation and activation function, respectively. Subsequently, we aggregate features from different stages through simple summation to obtain a comprehensive semantic representation that fuses both shallow and deep features:
$\mathbf{X}^{\text{sum}}$ $=\sum_{i}^{L}{\mathbf{X}^{i}}^{\prime}$ (3)
[…]
where $L$ denotes the number of RHAG blocks in the backbone. The resulting $\mathbf{X}^{\text{agg}}$ serves as the MFA output and will be used as input to the router for activating different experts in the upsampling process.
II-C Dual-Routing Heterogeneous Mixture of Experts
We propose the Dual-Routing Heterogeneous Mixture of Experts (DR-HMoE) to enhance the upsampling process in super-resolution networks. The DR-HMoE architecture comprises heterogeneous expert networks and their corresponding routers.
II-C1 Heterogeneous Experts
Given a set of expert networks $E_{ij}$, where $i\in\{1,2,\cdots,N\}$ and $j\in\{1,2,\cdots,M\}$, each $E_{ij}$ represents the $j$-th expert in the $i$-th expert group. Each expert network comprises a convolutional layer followed by a pixel shuffle operation, which can be formally expressed as:
[…]
where Conv2D denotes the convolutional operation and PS represents the pixel shuffle upsampling operation. To accommodate diverse types and scales of ground objects, experts across different groups utilize convolution kernels of varying sizes. While experts within each group share identical architectural structures, their parameters converge to distinct points in the parameter space during training. This heterogeneous architecture enables adaptive upsampling by providing specialized reconstruction functions for different pixels, thereby effectively handling various ground object patterns.
II-C2 Dual-Routing Mechanism
We propose a dual-routing strategy where pixel-level features are first directed to different expert groups and subsequently routed to the optimal experts within the selected group. The expert group router processes the aggregated features $\mathbf{X}^{\text{agg}}$ as input, as detailed in Section II-B. For each pixel-level feature vector $\mathbf{X}^{\text{agg}}_{k}$, the selection probability for each expert group is computed through a linear transformation followed by a softmax operation:
$p_{i}^{k}=\frac{e^{\mathbf{W}_{i}\cdot\mathbf{x}^{\text{agg}}_{k}}}{\sum_{i=1}^{N}e^{\mathbf{W}_{i}\cdot\mathbf{x}^{\text{agg}}_{k}}}$ (5)
where $\mathbf{W}_{i}\in\mathbb{R}^{C\times N}$ represents the weights of the linear transformation for the $i$th expert group, $i\in\{1,2,\cdots,N\}$, $k$ denotes the $k$-th feature vector in the aggregated feature map, and $N$ represents the total number of expert groups.
Following the computation of selection probabilities, $\mathbf{X}^{\text{agg}}_{k}$ is routed to the expert group with the highest probability, where experts within that group are activated for upsampling. The activation probabilities of individual experts are computed based on the aggregated features and position encoding of the expert group index:
$p_{j}^{k}=\frac{e^{\mathbf{W}^{i}_{j}\cdot(\mathbf{x}^{\text{agg}}_{k}+\text{PE})}}{\sum_{j=1}^{M}e^{\mathbf{W}^{i}_{j}\cdot(\mathbf{x}^{\text{agg}}_{k}+\text{PE})}}$ (6)
where $j\in\{1,2,\cdots,M\}$ represents the $j$-th expert in the activated expert group $i$, $\mathbf{W}^{i}_{j}\in\mathbb{R}^{C\times M}$ represents the weights of the linear transformation for the $j$th expert in $i$-th group, and PE denotes the position encoding of the activated expert group. Based on the activation probabilities $p_{j}^{k}$, the top-K experts (denoted as $\Omega$) are selected to process the low-resolution features ($\mathbf{x}^{\text{feat}}_{k}$) according to:
$\mathbf{X}_{\text{out}}^{k}=\sum_{j\in\Omega}p_{j}^{k}(\mathbf{x}^{agg}_{k})E_{j}(\mathbf{x}^{\text{feat}}_{k})$ (7)
where $\mathbf{X}_{\text{out}}^{k}$ represents the output pixel value, and $\Omega$ denotes the indices of the selected experts. Through this approach, each pixel is adaptively routed to an expert group and its most suitable experts for upsampling and reconstruction.
TABLE I: Comparison of different methods on UCMerced dataset.
Method ×$\times 2$$\times 4$
SSIM PSNR SSIM PSNR
EDSR [12] 0.9383 35.5104 0.7840 28.9274
RRDBNet [16] 0.9335 35.0327 0.7839 28.9929
RCAN [13] 0.9386 35.5456 0.7860 29.0668
SwinIR [17] 0.9390 35.5463 0.7877 29.0573
HAT [9] 0.9398 35.6829 0.7893 29.0985
HAUNet [18] 0.9389 35.6016 0.7899 29.0804
TTST [10] 0.9400 35.7058 0.7896 29.1455
MFG-HMoE (Ours) 0.9409 35.8110 0.7954 29.2882
TABLE II: Comparison of different methods on AID dataset.
Method ×$\times 2$$\times 4$
SSIM PSNR SSIM PSNR
EDSR [12] 0.9438 36.4926 0.8091 30.5434
RRDBNet [16] 0.9407 36.2417 0.8063 30.4646
RCAN [13] 0.9441 36.5261 0.8098 30.5880
SwinIR [17] 0.9434 36.4428 0.8089 30.5357
HAT [9] 0.9439 36.4975 0.8104 30.5920
HAUNet [18] 0.9436 36.4883 0.8079 30.5571
TTST [10] 0.9441 36.5142 0.8108 30.6108
MFG-HMoE (Ours) 0.9443 36.5399 0.8110 30.6167
TABLE III: Ablation studies in $\times$ 4 SR task on the UCMerced Dataset.
ID MFA Dual Routing Expert Num $1\times 1$ Experts $3\times 3$ Experts $5\times 5$ Experts PSNR SSIM
1 1 0 1 0 0.7884 29.1013
16 0 16 0 0.7924 29.1938
2 16 0 16 0 0.7924 29.1938
$\checkmark$ 16 0 16 0 0.7934 29.2631
3 $\checkmark$ 8 0 8 0 0.7932 29.2578
$\checkmark$ 16 0 16 0 0.7934 29.2631
$\checkmark$ 32 0 32 0 0.7939 29.2136
4 $\checkmark$ 16 16 0 0 0.7346 27.7779
$\checkmark$ 16 0 16 0 0.7934 29.2631
$\checkmark$ 16 0 0 16 0.7415 27.9768
5 $\checkmark$ $\checkmark$ 16 8 8 0 0.7954 29.2882
$\checkmark$ $\checkmark$ 16 0 8 8 0.7931 29.2075
$\checkmark$ $\checkmark$ 16 5 6 5 0.7937 29.2421
III Experiments
III-A Datasets and Experimental Setup
In this study, we evaluate the effectiveness of our proposed method using two datasets: UCMerced [14] and AID [15].
The UCMerced dataset consists of remote sensing images encompassing 21 distinct scene categories, with 100 images per category. Each image has dimensions of $256\times 256$ pixels and a spatial resolution of 0.3 meters per pixel. We randomly partitioned the images within each scene category into training and testing sets using a 3:1 ratio, yielding 1,575 training images and 525 testing images.
The AID dataset contains 10,000 remote sensing images distributed across 30 categories. Each image measures $600\times 600$ pixels with a spatial resolution of 0.5 meters per pixel. Similarly, we randomly split the images into training and testing sets using a 4:1 ratio, resulting in 8,000 training images and 2,000 testing images.
For both datasets, we used the original images as high-resolution (HR) ground truth and generated the corresponding low-resolution (LR) images through bilinear interpolation at specific downsampling scales. We compared our approach with several state-of-the-art methods, including EDSR [12], RRDBNet [16], RCAN [13], SwinIR [17], HAT [9], HAUNet [18], and TTST [10].
To ensure fair comparison, all experiments were conducted using an NVIDIA GeForce RTX 4090 GPU. Each model was trained for 100,000 iterations, and the best-performing model was selected for evaluation. We employed SSIM and PSNR as performance metrics. In the proposed model, we set the parameters as follows: $N=2$, $M=8$, and $K=1$.
III-B Comparison with Other Methods
We performed both qualitative and quantitative evaluations to compare our method with other state-of-the-art approaches. Figure 2 illustrates the ×4 super-resolution (SR) results on the UCMerced dataset. The comparison reveals that existing methods encounter difficulties in accurately reconstructing critical geospatial features. Specifically, EDSR, TTST, and SwinIR exhibit a tendency to generate non-existent lines, while HAT produces blurred line features. In contrast, our method demonstrates superior performance in reconstructing geospatial details with clarity and accuracy.
Tables I and II present the quantitative performance metrics for ×2 and ×4 SR tasks on both UCMerced and AID datasets. The results in Table I indicate that our method achieves substantial improvements over existing state-of-the-art approaches on the UCMerced dataset, particularly in the more challenging ×4 SR task. Furthermore, Table II shows that our method outperforms HAT in both SSIM and PSNR metrics without requiring modifications to the feature extraction network. Notably, these performance improvements exceed those achieved by TTST, despite TTST having a stronger feature extraction network.
Figure 2: Comparison of results from different methods in $\times$4 SR on the UCMerced dataset with the HR ground truth.
Figure 3: Visualization of the selected experts in $\times$4 SR on the UCMerced dataset.
III-C Ablation Studies
Table III presents the ablation study results on the UCMerced dataset. The first row shows the baseline model with a single upsampling layer. The experimental results demonstrate several key findings: 1) Group 1 reveals the impact of MoE that incorporating 16 experts significantly improves the super-resolution performance. 2) Group 2 shows that adding MFA while maintaining 16 experts leads to further performance gains by effectively leveraging both deep and shallow features, thus maximizing the potential of the MoE model for super-resolution tasks. 3) Group 3 compares different numbers of experts, indicating that optimal performance requires an appropriate expert count, which depends on the dataset characteristics and feature distribution. 4) Group 4 demonstrates the crucial role of convolution sizes in upsampling, with $3\times 3$ kernels achieving optimal results. 5) Group 5 shows that introducing a dual routing mechanism further enhances super-resolution capability. While keeping the total number of experts constant, experiments with varying numbers and types of heterogeneous expert groups reveal that the combination of $1\times 1$ and $3\times 3$ heterogeneous experts yields the best performance.
Additionally, we visualize the upsampling experts corresponding to each pixel, as shown in Fig. 3. The results indicate that for relatively smooth regions, such as vegetation, water bodies, and shadows, or for small-scale ground objects like vehicles, the model tends to select experts from Expert Group 1 for processing. In contrast, for larger-scale ground objects with richer details, such as buildings and oil tanks, the model favors selecting experts from Expert Group 2. This finding confirms that the model assigns different types of ground objects to different experts for processing, further validating the necessity of designing heterogeneous experts.
IV Conclusion
This paper presents a novel Multi-level Feature-Guided Heterogeneous Mixture of Experts (MFG-HMoE) model for remote sensing image super-resolution reconstruction. This model addresses the limitation of existing methods in distinguishing and reconstructing different ground object details. Our proposed framework employs multiple expert groups that maintain a balance between intra-group homogeneity and inter-group heterogeneity, facilitating specialized parameter activation for processing diverse ground details. Specifically, we introduce a Multi-level Feature Aggregation (MFA) module, which integrates deep and shallow features from the feature extraction network while providing crucial routing gate information for the Mixture of Experts (MoE). We propose a dual routing mechanism that optimizes upsampling expert assignment for individual pixels through a two-stage process. Experimental results on the UCMerced and AID datasets demonstrate that MFG-HMoE significantly surpasses state-of-the-art methods in super-resolution reconstruction accuracy. Through comprehensive ablation studies, we validate the effectiveness of each model component and establish optimal parameter configurations. This study advances the field of remote sensing image super-resolution reconstruction by introducing an innovative approach with promising implications for future research.